OpenStudy (anonymous):

ive tried solving this a few times but I keep getting an answer that's not one of the options, help please :) solve for x: 2/(x+2) + (1/5) = 6/(x+5)

OpenStudy (mimi_x3):

\[\large \frac{2}{x+2} +\frac{1}{5} = \frac{6}{x+5} \]

OpenStudy (anonymous):

I have to solve for x

OpenStudy (mimi_x3):

\[\frac{2*5}{(x+2)*5} +\frac{1*(x+2}{5*(x+2)} = \frac{6}{x+5} => \frac{10}{5(x+2)} + \frac{(x+2)}{5(x+2)} = \frac{6}{x+5} \]

OpenStudy (mimi_x3):

\[\frac{10+(x+2)}{5(x+2)} = \frac{6}{x+5} \]

OpenStudy (anonymous):

im confused so does x= 0?

OpenStudy (mimi_x3):

There are two solutions towards this question..

OpenStudy (anonymous):

0 and 13?

OpenStudy (mimi_x3):

Yes!

OpenStudy (anonymous):

yay! okay thank you:)

OpenStudy (mimi_x3):

np